The length of the longer leg of a right triangle is
25 inches more than seven times the length of the shorter leg. The length of the hypotenuse is 26 inches more than seven times the length of the shorter leg. Find the side lengths of the triangle.

Answer by Boreal(15235)

short leg is x
longer leg is 7x+25
the hypotenuse is 7x+26
square everything
x^2+49x^2+350X+625=49X^2+364X+676
X^2-14x-51=0
(x-17)(x+3)=0
x=17 inches only positive root
longer leg is 144 in
hypotenuse is 145 in